# When Do LLMs Admit Their Mistakes? Understanding the Role of Model Belief in Retraction

### v2

@@ 4.2 Steering Internal Beliefs Affects Retraction
We still use the UTQA training set to find steering directions. For each layer $l\in|L|$, we calculate the mean hidden states $h^{+}_{l}$ for correct answers at the last token of the answer, and $h^{-}_{l}$ for incorrect answers. We then compute the different-in-means vector $v_{l}=h^{+}_{l}-h^{-}_{l}$ (Marks and Tegmark, 2023b; Arditi et al., 2024; Li et al., 2023), which represents a linear belief direction. We add or subtract this difference-in-means vector to the activations of a new answer, thereby shifting the model’s perception of the correctness of the answer: $h_{l}^{\prime}\leftarrow h_{l}+\alpha v_{l}$, where $\alpha$ represents the strength of steering. Note that we steer only at the last token of the answer; we do not add the steering vector at any following generation steps in order to minimize disruption to the model’s natural generation. Similar to prior work (Turner et al., 2023; Lee et al., 2025; Li et al., 2023), we manually search for the steering hyperparameters to ensure that the steering is effective and minimally invasive, as detailed in Appendix B.4.
Figure 3: Retraction Rate under belief steering.
@@ Results.
We present retraction rate (i.e., the proportion of retracted examples) in Figure 3 for clarity and provide detailed retraction recall and precision in Appendix Table 12,  14, and 14. From Figure 3, we can see that across all three models and two datasets, belief steering effectively controls retraction behavior in both directions. Specifically, strengthening the model’s belief in the negative direction causes it to retract over 70% of the time across the entire dataset. In contrast, when we strengthen the model’s belief in the positive direction, retraction rate drops to nearly zero, indicating the model rarely retracts. This supports our hypothesis about the role of model belief in retraction: an LLM tends to take back an answer only when it internally believes it is incorrect; otherwise, it is like to stand by its initial answer.
We note that other steering directions, e.g., ones directly derived from in-distribution data, can yield similar results as detailed in Appendix C.2. However, these directions often fail to generalize to out-of-distribution settings. Importantly, our goal is not to find the optimal steering direction. Instead, we aim to understand when and why LLMs choose to retract. Both the probing and steering results support the conclusion that the model’s belief—defined independently of retraction and trained on separate data—causally affects retraction behavior and generalizes across different datasets.
5 Mechanistic Analysis
Having established that retraction behavior is guided by LLMs’ internal beliefs, we now turn to a deeper investigation of how beliefs function. In this section, we explore the mechanisms through which beliefs shape model behavior, from surface-level token generation to deeper attention dynamics.
5.1 Internal Beliefs Influence the Decision to Stop Generating
First, we find that belief steering controls whether the model stops generation immediately after the given answer. If the model outputs a “.” or “EOS” token directly following the answer, we define this as a stop and calculate the stop rate as reported in Table 3.
@@ 5.1 Internal Beliefs Influence the Decision to Stop Generating
Table 3: Stop Rate, which refers to the proportion of examples where the model stops generating right after the given answer.
We observe that positive belief steering increases stop rate, suggesting that when the model believes the answer is true, it is more likely to terminate generation early, foregoing the opportunity to verify the answer. In contrast, negative belief steering reduces stop rate: the model tends to generate additional information like the entity’s birthplace and profession, which encourages it to reflect on and potentially challenge its initial answer, even if ultimately retraction does not occur.
At the same time, belief steering does more than just changing immediate next token, as evidenced by the stop rate of Qwen2.5-7B and Olmo2-7B. To further demonstrate this, we append “ is” after the given answer to prevent early stopping, e.g., “Hillary Clinton is[Model generation continues from here…]”. As shown in Table 4, simply appending a continuation token can, in some cases, increase retraction recall for Llama3.1-8B, leading to improved retraction performance. Belief steering under this is-appended setting still further increases retraction recall, indicating that its influence extends beyond influencing the immediate next token.
@@
Precision Recall Precision Recall
@@
5.2 Beliefs Influence Retraction at Later Tokens Primarily via Attention Value Vectors
So far, we have shown that belief steering influences retraction behavior after the token following the answer. Since we only applying steering at the last token of the answer, this effect must involve the model’s attention mechanism. Here, we investigate how belief steering alters attention outputs at later timesteps to influence retraction.
Belief steering changes attention weights.
We start by measuring how belief steering changes attention weights. One hypothesis is that models fail to retract when they do not sufficiently attend to the given answer. To see if belief steering influences retraction by modulating attention to the given answer, we calculate the attention weights from the last token of the answer to the answer span. Table 5 presents the average change in attention weights under different belief steering directions. Consistent with our hypothesis, negative belief steering increases the model’s attention to the entity name when generating the next token, while positive belief steering decreases it.
Llama3.1-8B Qwen2.5-7B Olmo2-7B
Wiki. Celeb. Wiki. Celeb. Wiki. Celeb.
[…]
Table 5: Change in attention weights to the answer span.
@@ 5.2 Beliefs Influence Retraction at Later Tokens Primarily via Attention Value Vectors
Is this change in attention weights the primary way that beliefs influence retraction? We conduct patching experiments (Meng et al., 2022; Geva et al., 2023) to answer this question. Instead of directly adding steering vectors to the hidden states of each layer, we selectively retain specific components, such as attention weights or attention value vectors, from the steered model, and patch them into an unsteered model. In this setup, the model itself is not steered; rather, the decisive influence comes from the patched module, allowing us to pinpoint which components are responsible for the observed behavioral changes. We experiment with patching attention weights from salient heads (i.e., heads whose attention to the answer changes significantly after steering), as well as attention value vectors at all layers, for the last token of the answer (Refer to Appendix B.5 for implementation details). We present patching results for Llama3.1-8B in Table 7.
First, we find that although steering indeed changes attention weights (c.f., Table 5), patching attention weights alone has a relatively minor impact on retraction recall, especially under negative steering. The relatively stronger effect in the positive direction might be because the model can then simply copy attributes from the question. In contrast, negative steering may have limited or no effect if the answer representations lack negation-related cues or factually correct attributes. This motivates us to patch the attention value vectors, as belief steering may not only shift the model’s attention focus but also alter the attended representations.
Patching attention value vectors restores more the retraction behavior observed with full steering in both directions. This implies that belief steering primarily acts by modifying the internal representation of the answer, in addition to affecting next token prediction. In Table 7, we also present patching results for Llama3.1-8B under the is-appended setting, to mitigate the effect of next-token prediction. When this influence is reduced, attention value vectors play a more prominent role. This is also verified by experiments on Qwen2.5-7B and Olmo2-7B as shown in Appendix C.3.
Wikidata Celebrity
Prec. Rec. Prec. Rec.
No Steer 0.9012 0.2579 0.7722 0.1477
belief-
Patch W 0.8325 0.2729 0.7113 0.1671
Patch V 0.5249 0.5441 0.6351 0.3245
Full Steer 0.5157 0.9268 0.4803 0.7676
belief+
Patch W 0.8984 0.1913 0.7333 0.1065
Patch V 0.9700 0.1614 0.6552 0.0920
Full Steer 1.0000 0.0067 0.5217 0.0291
Table 6: Patching results for Llama3.1-8B on continuation test sets.
@@ 6 Supervised Fine-Tuning Can Learn Better Internal Belief
Given that supervised fine-tuning can enhance existing capabilities of LLMs (Prakash et al., 2024; Yang et al., 2024c), do our findings on the role of model beliefs in retraction generalize to fine-tuning models with enhanced retraction performance? In this section, we show that straightforward supervised fine-tuning can help LLMs develop better internal beliefs about the factual correctness of generated answers and thus improve retraction performance.
@@
Precision Recall Precision Recall
Baseline 0.9012 0.2529 0.7722 0.1477
@@
For each of our datasets, we synthetically construct an in-distribution supervised fine-tuning training set (e.g., using the Wikidata training set). Specifically, we append “ is the correct answer.” to correct examples that contain factually correct answers in the training dataset, and “ is not the correct answer.” to wrong examples, and use LoRA (Hu et al., 2022) to fine-tune LLMs. The results of Llama3.1-8B are shown in Table 8. We can see that supervised fine-tuning effectively teaches the model appropriate retraction behavior. The model learns to distinguish between factually correct and incorrect answers and respond accordingly, i.e., saying “is the correct answer” to correct answers and saying “is not the correct answer” to incorrect ones. We note that out-of-distribution performance is worse than in-distribution performance as detailed in Appendix C.4.1; we view our findings as a proof-of-concept and hypothesize that a larger and more diverse training dataset could yield robust retraction capabilities.
@@
Then we apply the same belief steering vectors from the original model and the same hyperparameters [^2] Note that these may not be the optimal hyperparameters. In fact, extending steering from layers 6-14 to 6-20 reduces retraction recall on Belief+ Celebrity set from 0.5763 to 0.2300 with no change in response format. to steer the fine-tuned model. As shown in Table 8, the steering vectors can be generalized to the fine-tuned model and change its retraction behavior in both directions, without altering its response format, i.e., “is (not) the correct answer”. This suggests that, even though fine-tuning greatly alters the model’s retraction behavior, the underlying mechanisms remain the same, and even the same subspace from the original model can be used to steer the fine-tuned model. Similar results for Qwen2.5-7B and Olmo2-7B, presented in Appendix C.4.2, further confirm this observation.
(a) Llama3.1-8B on Wikidata.
(b) Llama3.1-8B on Celebrity.
Figure 4: Average probe scores across layers for Llama3.1-8B (Base) and its fine-tuned variant (SFT). “C” denotes correct examples, and “W” denotes wrong examples.
Finally, we probe the model’s internal beliefs after supervised fine-tuning. As shown in Figure 4, factually incorrect answers previously receive high probe prediction scores, indicating that the model tends to treat them as correct. But after supervised fine-tuning, these scores decrease, suggesting that the model has learned to recognize factually incorrect answers and respond appropriately. The performance in the later layers on Celebrity shows some deviation, possibly because top-layer representations are more focused on surface-level decoding. Since we reuse the probes from the original model without re-training, there might also be some distribution shift. Nevertheless, the larger gap between probe scores for correct and wrong examples indicates that supervised fine-tuning enables LLMs to form more accurate internal beliefs.
7 Conclusions and Limitations
@@
In this paper, we evaluate and analyze the underlying mechanisms behind retraction in LLMs. Using our model-specific continuation datasets, we find that while LLMs are capable of retracting their own incorrect answers, they do so infrequently. Through probing and steering experiments, we demonstrate that retraction is causally influenced by the model’s internal belief: a model fails to retract an incorrect answer because it genuinely believes it is correct. We further show that beliefs guide retraction by affecting both the surface-level token predictions and deeper attention dynamics. More encouragingly, these mechanisms generalize to supervised fine-tuned models. Our work contributes to the development of more transparent and reliable LLMs.
@@
There are several limitations for future research. First, although different LLMs share the same overall retraction mechanism—being causally influenced by the model’s internal belief—the specific layers where this influence is most pronounced vary across models. As shown in Appendix B.4, steering at early to mid layers is effective for Llama3.1-8B and Qwen2.5-7B, whereas Olmo2-7B requires intervention at higher layers to elicit stronger retraction. These differences likely stem from variations in training recipes, including data and optimization strategies used.
@@
Second, our analysis focuses on short-answer knowledge-related question answering tasks. One natural extension is to long-form generation, such as “Name 15 politicians who were born in New York City”. This introduces new challenges, including how to accurately locate each generated answer and how to isolate the influence of earlier outputs on later ones. Moreover, as many self-correction studies target reasoning tasks, it would be valuable to examine whether our findings generalize to that domain. However, caution is needed to disentangle limitations in retraction from other capabilities required for reasoning tasks, such as arithmetic computation and problem understanding.
@@
Appendix A Preliminary Analysis of Large Reasoning Models
Recent large reasoning models (LRMs) such as OpenAI o1 (OpenAI, 2024), DeepSeek-R1 (DeepSeek-AI et al., 2025), and QwQ (Qwen Team, 2025a) are known for automatic self-reflection, where the model reflects on its own responses without any external hints or instructions in their thinking mode. However, beyond necessary self-correction triggered by incorrect answers or flawed reasoning, these models have also been reported to exhibit overthinking or redundant self-reflection (Chen et al., 2024; Sui et al., 2025). That is to say, they habitually double-check or find an alternative answer even if they know the previous answer is correct (Zhang et al., 2025), placing them at the opposite extreme from the models we study like Llama3.1-8B-Instruct.
Nonetheless, there appear to be connections between LLMs and LRMs. According to our experiments in Section 5.1, when Llama3.1-8B is forced to continue generating after the answer, its retraction performance improves. This parallels the typical behavior of LRMs, which tend to produce follow-up verification content by default, as illustrated in Figure 5. Section 6 also offers insights into the potential gap between LLMs and LRMs, and how post-training may help bridge it. Lastly, we observe that an LRM in non-thinking mode tends to behave like a non-thinking model (See Figure 6). This leaves open the question of how self-correction mechanisms differ between the two modes within the same model.
Figure 5: Qwen3-32B (Qwen Team, 2025b) in thinking mode.
Figure 6: Qwen3-32B in non-thinking mode.
Appendix B Experimental Details
B.1 Details of Original Datasets
@@
We focus on knowledge-related question answer tasks, where it is transparent whether an LLM has the necessary knowledge to identify its mistakes. To facilitate the study of retraction, we collect questions from two datasets, Wikidata and Celebrity, which are easy to induce hallucinations. The number of questions in each split of the datasets is reported in Table 9
@@
Wikidata was originally proposed by Dhuliawala et al. (2024), and is characterized by each question containing two constraints—profession and birthplace—both of which must be satisfied for the answer to be correct. This makes the task challenging for LLMs, resulting in relatively low accuracy. However, the original dataset was not publicly released. To reconstruct it, we collect a set of popular professions and cities, and generate new questions by pairing them. We retain only those combinations for which a correct exists. For accuracy evaluation, we query the Wikidata API [^3] https://query.Wikidata.org/. An example question is:
@@
B.2 Details of Continuation Datasets

### v3

@@ 4.2 Steering Internal Belief Affects Retraction
We still use the UTQA dataset to find steering directions. For each layer $l\in|L|$, we calculate the mean hidden states $h^{+}_{l}$ for correct answers at the last token of the answer, and $h^{-}_{l}$ for incorrect answers. We then compute the difference-in-means vector $v_{l}=h^{+}_{l}-h^{-}_{l}$ (Marks and Tegmark, 2023a; Arditi et al., 2024; Li et al., 2023), which represents a linear belief direction. We add or subtract this difference-in-means vector to the activations of a new answer, thereby shifting the model’s perception of the correctness of the answer: $h_{l}^{\prime}\leftarrow h_{l}\pm\alpha v_{l}$, where $\alpha$ controls the strength of steering. Note that we steer only at the last token of the answer; we do not add the steering vector at any following generation steps in order to minimize disruption to the model’s natural generation. Similar to prior work (Turner et al., 2023; Lee et al., 2025; Li et al., 2023), we manually search for the steering hyperparameters to ensure that the steering is effective and minimally invasive, as detailed in Appendix A.3.
Figure 3: Retraction rate under belief steering. “Belief-” denotes negative belief steering while “Belief+” denotes positive belief steering.
@@ Results.
We present retraction rate (i.e., the proportion of retracted examples) in Figure 3 for clarity and provide detailed retraction recall and precision in Appendix Table 13,  15, and 15. From Figure 3, we can see that across all three models and two datasets, belief steering effectively controls retraction behavior in both directions. Specifically, strengthening the model’s belief in the negative direction causes it to retract over 70% of the time across the entire dataset. In contrast, when we strengthen the model’s belief in the positive direction, retraction rate drops to nearly zero, indicating the model rarely retracts. This supports our hypothesis about the role of model belief in retraction: an LLM tends to take back an answer only when it internally believes it is incorrect; otherwise, it is likely to stand by its initial answer.
We note that the belief steering direction is derived from out-of-distribution data, yet remains effective at modulating retraction behavior across different datasets and under variations in prompting and decoding hyperparameters (Appendix B.3). This contrasts with common steering approaches, which typically rely on in-distribution data (Appendix B.4). More encouragingly, the causal link between belief and retraction extends beyond knowledge-based question answering. In GSM8k (Cobbe et al., 2021), a widely used math reasoning dataset, we identify cases where Llama3.1-8B makes an intermediate mistake. By applying negative belief steering at the token where the mistake happens, we induce the model to retract, leading it to ultimately get the correct answer in 20% of these cases (Appendix B.8).
Overall, we aim to understand when and why LLMs choose to retract. Taken together, both the probing and steering results support the conclusion that the model’s belief—defined independently of retraction and trained on separate data—causally affects retraction behavior and generalizes across different datasets.
@@
Table 3: Stop rate: the proportion of examples where the model stops generating right after the given answer.
5 Mechanistic Analysis
Having established that retraction is guided by LLMs’ internal beliefs, we now turn to a deeper investigation of how beliefs function. In this section, we explore the mechanisms through which beliefs shape model behavior, from surface-level token generation to deeper attention dynamics.
5.1 Belief Influences the Decision to Stop Generating
First, we find that belief steering controls whether the model stops generation immediately after the given answer. If the model outputs a “.” or “EOS” token directly following the answer, we define this as a stop and calculate the stop rate in Table 3.
We observe that positive belief steering increases stop rate, suggesting that when the model believes the answer is true, it is more likely to terminate generation early, foregoing the opportunity to verify the answer. In contrast, negative belief steering reduces stop rate: the model tends to generate additional information like the entity’s birthplace and profession, which encourages it to reflect on and potentially challenge its initial answer.
At the same time, belief steering does more than just change the immediate next token. To demonstrate this, we append “is” after the given answer to prevent early stopping, e.g., “Hillary Clinton is[Model generation continues from here…]”. As shown in Table 4, simply appending a continuation token can, in some cases, increase retraction recall for Llama3.1-8B, leading to improved retraction performance. Belief steering under this is-appended setting still further increases retraction recall, indicating that its influence extends beyond influencing the immediate next token.
@@ 5.1 Belief Influences the Decision to Stop Generating
Prec. Rec. Prec. Rec.
@@
5.2 Belief Influences Retraction Primarily via Attention Value Vectors
Given that retraction behavior is closely tied to previous context and involves the model’s attention mechanisms, we next examine how belief steering modifies attention outputs to influence retraction.
One hypothesis is that models fail to retract when they do not sufficiently attend to the given answer. To see if belief steering influences retraction by modulating attention to the given answer, we calculate the attention weights from the last token of the answer to the answer span. In Appendix, Table 16 presents the average change in attention weights under different belief steering directions. Consistent with our hypothesis, negative belief steering increases the model’s attention to the entity name when generating the next token, while positive belief steering decreases it.
@@ 5.2 Belief Influences Retraction Primarily via Attention Value Vectors
Is this change in attention weights the primary way that beliefs influence retraction? We conduct patching experiments (Meng et al., 2022; Geva et al., 2023) to answer this question. Instead of directly adding steering vectors to the hidden states of each layer, we selectively retain specific components, such as attention weights or attention value vectors, from the steered model, and patch them into an unsteered model. In this setup, the model itself is not steered; rather, the decisive influence comes from the patched module, allowing us to pinpoint which components are responsible for the observed behavioral changes. We experiment with patching attention weights from salient heads (i.e., heads whose attention to the answer changes significantly after steering), as well as attention value vectors at all layers, for the last token of the answer (Refer to Appendix A.4 for implementation details). We present patching results for Llama3.1-8B under the is-appended setting in Table 5, to mitigate the effect of next-token prediction.
@@
We can observe that patching attention weights (i.e., Patching W) does have a measurable effect on retraction; however, this effect is relatively modest. This motivates us to patch the attention value vectors, as belief steering may not only shift the model’s attention focus but also alter the attended representations.
(a) Llama3.1-8B on Wikidata.
(b) Llama3.1-8B on Celebrity.
Figure 4: Layer-wise AUROC of belief scores in Llama3.1-8B (Base) and its fine-tuned variant (SFT).
Compared to attention weights, patching attention value vectors (i.e., Patching V) restores more the retraction behavior observed with full steering (i.e., Full Steer) in both directions. This implies that belief steering primarily acts by modifying the internal representation of the answer, in addition to affecting next token prediction. This is further supported by experiments without appending is, as well as by consistent results on Qwen2.5-7B and Olmo2-7B as shown in Appendix B.6.
@@ 6 Supervised Fine-Tuning Can Learn Better Internal Belief
Given that SFT can enhance existing capabilities of LLMs (Prakash et al., 2024; Yang et al., 2024b), we first verify that straightforward SFT indeed improves in-distribution retraction performance. Specifically, we append “is the correct answer.” to correct examples and “is not the correct answer.” to wrong examples in the training data, thereby explicitly encouraging appropriate retraction behavior (training details can be found in Appendix A.5). As shown in Table 6, the fine-tuned model learns to distinguish factually correct from incorrect answers and to respond accordingly. Having established this, what remains underexplored is whether our findings on the role of model belief in retraction continue to hold after fine-tuning.
@@
Prec. Rec. Prec. Rec.
Baseline 0.9012 0.2579 0.7722 0.1477
@@
We first apply the same belief steering vectors from the original model and the same hyperparameters [^3] Note that these may not be the optimal hyperparameters. In fact, extending steering from layers 6-14 to 6-20 reduces retraction recall on Belief+ Celebrity set from 0.5763 to 0.2300. to steer the fine-tuned model. As shown in Table 6, the steering vectors can be generalized to the fine-tuned model and change its retraction behavior in both directions, without altering its response format learned during SFT, i.e., “is (not) the correct answer”. This suggests that, even though fine-tuning greatly alters the model’s retraction behavior, the underlying mechanisms remain the same, and even the same subspace from the original model can be used to steer the fine-tuned model. Similar results for Qwen2.5-7B and Olmo2-7B, presented in Appendix B.7, further confirm this observation.
We also probe the model’s internal belief after SFT, as shown in Figure 4. Since we reuse the probes from the original model without re-training, there might be some distribution shift. Nevertheless, these probes based on internal representations remain effective at predicting external retraction behavior. More interestingly, SFT aligns the model’s internal belief more closely with factual correctness, as reflected by the higher AUROC for Factual Correctness (SFT) compared to Factual Correctness (Base). This suggests that SFT enables LLMs to form more accurate internal beliefs.
7 Conclusions
@@
In this paper, we evaluate and analyze the underlying mechanisms behind retraction in LLMs. Using our model-specific continuation datasets, we find that while current LLMs are capable of retracting their own incorrect answers, they do so infrequently. Through probing and steering experiments, we demonstrate that retraction is causally influenced by the model’s internal belief: a model fails to retract an incorrect answer when it internally believes it is correct. We further show that beliefs guide retraction by affecting both the surface-level token predictions and deeper attention dynamics. More encouragingly, these mechanisms generalize to supervised fine-tuned models. We hope our work contributes to the development of more transparent and reliable LLMs.
Limitations
@@
There are several limitations for future research. First, although different LLMs share the same overall retraction mechanism—being causally influenced by the model’s internal belief—the specific layers where this influence is most pronounced vary across models. As shown in Appendix A.3, steering at early to mid layers is effective for Llama3.1-8B and Qwen2.5-7B, whereas Olmo2-7B requires intervention at higher layers to elicit stronger retraction. These differences likely stem from variations in training recipes, including data and optimization strategies used.
@@
Second, our analysis primarily focuses on short-answer, knowledge-based question answering tasks. One natural extension is to long-form generation, such as “Name 15 politicians who were born in New York City”. As many self-correction studies target reasoning tasks, it would also be valuable to examine whether our findings generalize to that domain. While we demonstrate that belief steering generalizes to GSM8k in preliminary experiments (see Appendix B.8), several challenges remain, including accurately identifying steering points and isolating the influence of earlier outputs on later ones.
Ethical Considerations
This work does not raise specific ethical concerns. All datasets used in this study (Wikidata, Celebrity, and UTQA) are publicly available and do not contain private or sensitive information. Our analysis focuses on model behavior and does not involve human subjects.
@@
Appendix A Experimental Details
A.1 Details of Original Datasets
@@
We focus on knowledge-related question answer tasks, where it is transparent whether an LLM has the necessary knowledge to identify its mistakes. To facilitate the study of retraction, we collect questions from Wikidata and Celebrity, which are easy to induce hallucinations. The number of questions in each split of the datasets is reported in Table 7
Question Properly Steered Response Oversteered Response
@@
Wikidata was originally proposed by Dhuliawala et al. (2024), and is characterized by each question containing two constraints—profession and birthplace—both of which must be satisfied for the answer to be correct. This makes the task challenging for LLMs, resulting in relatively low accuracy (Yüksekgönül et al., 2024). However, the original dataset was not publicly released. To reconstruct it, we collect a set of popular professions and cities, and generate new questions by pairing them. We retain only those combinations for which a correct answer exists. For accuracy evaluation, we query the Wikidata API [^4] https://query.Wikidata.org/. An example question is:
@@
A.2 Details of Continuation Datasets